OpenStudy (anonymous):

Round to the nearest tenth. The value of z is the value of a2 − 3a when a = -5. Find the value of z2.

OpenStudy (anonymous):

U just posted this... Clarify the question, it is unclear.

OpenStudy (anonymous):

this is what it says in my book

OpenStudy (anonymous):

It say a2 or a with a little 2?

OpenStudy (anonymous):

little 2

OpenStudy (saifoo.khan):

there is no z with the equation.

OpenStudy (anonymous):

Read questio, saifoo..

OpenStudy (anonymous):

my book must be cookoo

OpenStudy (saifoo.khan):

Lol, now i got it.

OpenStudy (anonymous):

a^2 = -5*-5 = 25 -3a = -3*-5 =15 a2-3a = 40 =z z^2 = z*z = 1600

OpenStudy (saifoo.khan):

i was correct Duck.

OpenStudy (saifoo.khan):

z = a2 − 3a

OpenStudy (saifoo.khan):

z = a2 − 3a z = 25 - 15 z = 10

OpenStudy (anonymous):

saifoo, a is negative....

OpenStudy (saifoo.khan):

Lol,

OpenStudy (saifoo.khan):

then the answer is: z = a2 − 3a z = 25 -3(-5) z = 25 + 15 z = 40

OpenStudy (anonymous):

and z^2 is..?

OpenStudy (saifoo.khan):

z^2 = 1600.. lol..

OpenStudy (anonymous):

You frightened him away...